The joint probability distribution of X and Y is given by

f(x,y)=x+y, 0<x<1 and 0<y<1 zero otherwise.

Determine the joint distribution function of X and Y i.e F(x,y). Hence or otherwise calculate p(0<x<1/2,1/2<y<1)


1
Expert's answer
2021-12-24T08:18:35-0500

joint distribution function:

"F(x,y)=P(X\le x,Y\le y)"


for "x\le 0" or "y\le 0" :

"F(x,y)=0"

for "x\ge 1" and "y\ge 1" :

"F(x,y)=1"


for "0<x<1 , 0<y<1" :


"F(x,y)=\int^y_0\int^x_0f(u,v)dudv=\int^y_0\int^x_0(u+v)dudv="


"=\int^y_0(x^2/2+vx)dv=yx^2/2+y^2x/2"


for "0<x<1" and "x\ge 1" :

"F(x,y)=F(x,1)=x^2\/2+x\/2"

for "0<y<1" and "y\ge 1" :

"F(x,y)=F(1,y)=y^2\/2+y\/2"


"P(0<x<1/2,1/2<y<1)=\int^1_{1/2}\int^{1/2}_0f(x,y)dxdy="


"=yx^2\/2+y^2x\/2|^{1\/2,1}_{0,1\/2}=1\/8+1\/2=5\/8"